# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 6.1.2. From a gash to a rigid lozenge
Let us first show a general fact about a shape described below that we call a cup. For $r\geq 1$, we call a cup of length $r$ the union of $r$ consecutive type 00 edges together with one edge of type $2$, respectively of type $1$, forming an angle of $\frac{2\pi}{3}$ with the west-most, respectively east-most type 00 edge. See Figure 16 for an example. Suppose that edges of a cup are all colored $1$. Let us show by induction on $r$ that the only possible color of edges in the convex hull of the cup is $1$.
Figure 16. A cup of length $r=8$. Edges in the convex hull of the cup are dotted
Two adjacent edges of different types form a type $1$ arrow. If this arrow had completion $(1.II)$ of Figure 13 all the $r$ edges of type 00 would belong in some lozenge with middle edge colored $3$ which is incompatible with the $1$ color of the type $1$ edge at the east of the cup. Thus, the arrow completion is $(1.I)$. Then, the rest of the convex hull is a cup of length $r-1$ which completes the proof of cup completion.
Considering $R$, the south-west corner triangular face has all of its edges colored $1$ as two of them lie on the boundary of $R$. North-east to this face, there is a cup of length $1$ having edges of type $1$. By the previous, its only completion is edges of color $1$. At each completion of a cup of size $r$, the north-east region is a cup of size $r+1$. Filling cups of sizes $11,2,\dots,s-$ 1 with edges of color $1$ fills $R$ proves the claim. ∎
By a similar argument, Lemma 6.9 also holds if two boundaries of $R$ are colored 00 by replacing $3$ lozenges by rigid lozenges in the proof.
Proof of Lemma 6.8.
Assume for the sake of contradiction that $g^{\prime}$ does not encounter any configuration $(v)$ or $(vi)$. By assumption, $g^{\prime}$ propagates via configurations $(i)$, $(ii)$ or $(iv)$ which preserve its type until reaching the boundary $\partial T_{n}^{(1)}$. Recall that $\partial T_{n}^{(1)}$ and $\partial T_{n}^{(0)}$ are regular. Therefore, the south east corner of $H$ consists of a $d$ equilateral triangle having all of its edges colored $1$ by Lemma 6.9. Let us call $R_{d}$ this region. Thus there cannot be a $(i)$ configuration on $\partial T_{n}^{(1)}$ so that the gash’s path cannot end by $(i)$. Let $l\geq 0$ be the number of configurations $(ii)$ from the last either $(i)$ or $(iv)$ configuration met. If this was configuration $(i)$, then we would have a 00 colored edge $e$ coming from the last configuration $(i)$ in $R_{d}$. Likewise, if it was configuration $(iv)$, consider the south-east most edge $e$ having color 00 in the last arrow of type $1$ with non-zero length (circled). Then, $e$ would be contained in $R_{d}$. Both cases would imply that $R_{d}$ has a 00 colored edge and would contradict the regularity assumption on $H$. See Figures 18 and 18.
Figure 17. The last $(i)$ configuration has a 00 colored edge inside $R_{d}$.
Figure 18. The last $(iv)$ configuration has a 00 colored edge inside $R_{d}$.
Proof of Proposition 6.7.
Assume that no configuration $(v)$ or $(vi)$ is encountered in step $(2^{\prime})$ when propagating a gash of type 00. Let us show that a configuration $(v)$ or $(vi)$ is encountered in the steps $(2)$ and $(3)$.
Consider a gash $g$ of type $1$ in $H$. The gash propagates a finite number of times via $(i)$, $(ii)$ and $(iv)$ configurations. As they both preserve the type of $g$, the propagation path must either end with $(v)$ or $(vi)$ or meet an upward $(iii)$ configuration $t$ from its west border. By Lemma 6.8, the gash $g^{\prime}$ on the east border of $t$ either meets a downward $(iii)$ or a configuration $(v)$ or $(vi)$. From a downward $(iii)$, $g^{\prime}$ propagates to a gash $g^{\prime\prime}$ of type $1$ associated to its east side. As there can only be a finite number of propagations $g\rightarrow g^{\prime}\rightarrow g^{\prime\prime}$. Consider the last $(iii)$ configuration $t$ met either upward or downward. If $t$ is upward, Lemma 6.8 shows that the path ends in a configuration $(v)$ or $(vi)$. Moreover, $t$ cannot be downward as the next gash in the propagation would have type $1$ which cannot reach $\partial T_{n}^{(1)}\cup\partial T_{n}^{(0)}$ due to its type preserved by configurations $(i)$, $(ii)$ and $(iv)$. Therefore, the path from $t$ ends in a configuration $(v)$ or $(vi)$. ∎
6.2. Hexagon rotation
Definition 6.10 (ABC hexagons).
Let $C$ be a color map and let $h$ be a hexagon, that is, the union of six triangular faces sharing one vertex in $T_{n}$. We say that $h$ is an ABC hexagon (for the color map $C$) if the color map $C$ restricted to $h$ is any of the three configurations in Figure 19 up to rotations which preserves the value of $C$ on the boundary $\partial h\cap E_{n}$. Note that type B has three possible rotations whereas A and C only have two. Denote $E_{h}$ the edges of $E_{n}$ which are in $h$.
A rotation of an ABC hexagon $h$ is the replacement of the values of $C$ by the ones obtained from a rotation of $h$ that preserve its boundary.
6.3. Color swap path
In Proposition 6.7, we showed that any gash $g$ on $\partial T_{n}^{(2)}$ can be propagated to find a configuration $(v)$ or $(vi)$ having a rigid lozenge in it. In this section, we show that via hexagon rotations of Definition 6.10, one can bring this rigid lozenge on $\partial T_{n}^{(2)}$ at the location of $g$. We first show that such a lozenge can travel through a configuration $(iv)$ in Lemma 6.11 below.
Lemma 6.11 (Color swap in region $(iv)$).
Let $g$ be a gash having a configuration $(iv)$ adjacent to it. Let $g^{\prime}$ be the propagation of $g$ given by Definition 6.6. Suppose that $g^{\prime}$ is adjacent to a configuration $c$ of type $(v)$ or $(vi)$. Using hexagon rotations, one can place $c$ so that $g$ is adjacent to $c$.
Proof.
We will do the proof in the case of configuration $(vi)$ but the same can be adapted in the case of configuration $(v)$ by changing the types of the hexagons moved in the process. Denote $R$ the region induced by the configuration $(iv)$. Suppose that $R$ consists only in the region delimited by two non-zero length arrows, one of type 00 and one of type $1$ as in Figure 14. Notice that using hexagon rotations (of type A if the arrow is type $1$ and type C if arrow is type 00) an arrow can be reversed i.e the resulting configuration is the arrow rotated by an angle $\pi$. We start by reversing the top type 00 arrow. Note that this creates an arrow of the same type below it. By induction, reverse all the $l_{1}$ arrows in the region. Note that the hexagon whose diagonal is $g^{\prime}$ is now of type B. Rotate it such that the rigid lozenge $l$ from the configuration $(vi)$ east to $g^{\prime}$ is vertical. Using type A rotations, one can bring $l$ to the south-west corner of $R$ in a type B hexagon. Doing a B hexagon rotation and then type C ones, one can place $l$ adjacent to $g$ forming a configuration (vi). The procedure is illustrated in Figures 21 to 25 below where for the sake of clarity we omit the $m$ colors from Figure 21, keeping track only of the rigid lozenge originally in $(vi)$ denoted $M$.
Figure 20. Original configuration delimited by two arrows.
Figure 21. Reverting type 00 arrows.
Figure 22. Type B hexagon rotation.
Figure 23. Type A hexagon rotations.
Figure 24. Type B hexagon rotation.
Figure 25. Type C hexagon rotations.
In the general case, by Lemma 6.5 the region $R$ consists of an amalgamation of regions $(l_{1},l_{2})$ similar to the one in Figure 21. Applying the previous procedure to each of those regions, one can place $c$ so that $g$ is adjacent to $c$. Notice that all the hexagons rotations in this proof keep the boundaries of the region fixed. ∎
Proposition 6.12 (Color swap path).
Let $g$ be a gash on $\partial T_{n}^{(2)}$. Consider the first $(v)$ or $(vi)$ configuration encountered in the algorithm of Proposition 6.7. This configuration has a rigid lozenge $c$ and let us call color swap path the resulting path between $g$ and $c$. Using hexagonal rotations on this path, one can move $c$ so that $g$ is adjacent to $c$.
Proof.
If a configuration of type $(v)$ or $(vi)$ is adjacent to the exit gash of a step $(i)$ or $(ii)$, the resulting hexagon is of ABC type and can be rotated so that $c$ is now on the entry gash. Lemma 6.11 shows that the same holds if $c$ is on the exit gash of a region $(iv)$. One can also check that if $c$ is on the bottom side or east side of a configuration $(iii)$, one can place it on the west border via a hexagon rotation. Each of those rotations keep boundaries of the path fixed. Therefore, one can apply hexagon rotations on the path from $g$ to $c$ which result in $g$ and $c$ being adjacent. ∎
6.4. Color map reduction
In this section, we only consider regular boundary conditions for $T_{n}$ that is, the color map $C:E_{n}\rightarrow\{0,1,3,m\}$ is given by $1\dots 10\dots 01\dots 1$ on every boundary of $T_{n}$, where there are $d$ ones one each side of the zeros. The goal of this section is to show that any color map having regular boundaries can be mapped via hexagon rotation to the simple color map of Definition 6.13.
Definition 6.13 (Simple color map).
Let $k,d\geq 0$ and $n=k+d$. The color map $C_{0}:E_{n}\rightarrow\{0,1,3,m\}$ called the simple color map is defined by
  1. (1)
    $C_{0}$ is given by $1\dots 10\dots 01\dots 1$ on every boundary of $E_{n}$, where there are $d$ ones around the zeros,
  2. (2)
    $C_{0}(e)=1$ for every edge $e\in E_{n}$ in any corner equilateral triangle of side length $d$ in $T_{n}$,
  3. (3)
    $C_{0}(e)=1$ for every edge $e\in E_{n}$ in the vertical lozenge of side length $d$ in $T_{n}$, two of its sides being the $d$ north-most colored edges of $\partial_{\lambda}T_{n}$ and $\partial_{\mu}T_{n}$,
  4. (4)
    $C_{0}(e)=0$ for every edge $e\in E_{n}$ in the equilateral triangle having one of its side consisting of the type 00 edges $\partial_{\nu}E_{n}$,
  5. (5)
    $C_{0}(e)=m$, respectively $C_{0}(e)=3$ for every horizontal edge $e\in E_{n}$ in the $(k-d)\times d$ rhombus in-between the triangle of $(4)$ and the 00 labels of $\partial_{\lambda}T_{n}$, respectively $\partial_{\mu}T_{n}$.
Here is an example of the simple color map $C_{0}$ for $n=5$ and $d=2$.
Figure 26. The color map $C_{0}$ for $n=5$ and $d=2$. Uncolored edges have color $3$.
Proposition 6.14 (Color map reduction).
Let $C:E_{n}\rightarrow\{0,1,3,m\}$ be a color map with regular boundaries. Using hexagon rotations, one can map $C$ to $C_{0}$ the simple color map of Definition 6.13.
Proof.
Let $C$ be a color map on $E_{n}$. Since $C$ is regular, the $d$ equilateral triangles on each corner of $T_{n}$ consists in triangular faces having boundary colors $C(e)=1$ by Lemma 6.9. Consider the south-west such corner. Its north-most $1$ colored edge of type $1$ forms with the south-most 00 colored edge on $\partial T_{n}^{(2)}$ an angle of $\frac{2\pi}{3}$ that we will call a $01$ opening. A $01$ opening has only three possible completions showed in Figure 27 depending on the color of the horizontal edge in the angle.
@@ 6.4. Color map reduction
If both the two edges belong to a same rigid lozenge as in the first completion in Figure 27, the south-east edge of type $1$ and color 00 of that lozenge forms a $01$ opening with the $1$ colored edge below it. If all these successive $01$ openings are part of rigid lozenge, then there are $d$ such rigid lozenges and the triangular face south to the last one has all of its edges colored 00 as two of its edges are. Then, the second south-most 00 label on $\partial T_{n}^{(2)}$ is part of a $01$ opening. Consider the first $01$ opening $o$ not filled with a rigid lozenge with respect to that exploration order. Since $o$ is not part of a rigid lozenge, the region east to $o$ is one of two rightmost possibilities of Figure 27. Both of them give rise to a $01$ gash $g$ on their east side. The region west to $o$ consists of vertical rigid lozenges on top of each other.
By Proposition 6.12 there exists a rigid lozenge in the region east to $g$ that can be moved to $g$ via hexagon rotations. Therefore, the $01$ opening $o$ can be filled using a rigid lozenge without changing the region west to it. Therefore, $C$ is equivalent up to hexagon rotations to the color map where the $(k-d)\times d$ rhombus east to $\partial T_{n}^{(2)}$ consists only of rigid lozenges. From this configuration, there is only one possible color map to complete the rest of the hive $E_{n}$ which is the simple color map $C_{0}$. ∎
Since the number of piece of each type is preserved under hexagon rotations, one derive the following enumerations.
@@
Let $H$ be a hive with regular boundaries. Let $hc(H)$, respectively, $sc(H),$ be the number of $m$, respectively $3$ colored edges inside $H$. Then,
@@
Moreover, for $i\in\{0,1\}$, denote $nt^{(i)}(H)$, respectively $st^{(i)}(H)$, the number of north-pointing, respectively south-pointing, triangular pieces of size $1$ with color $i$ on each side. Then,
@@
where $n(0)=n-d$ and $n(1)=2d$ which gives $nt^{(0)}(H)=\frac{(n-d)(n-d+1)}{2}$, $st^{(0)}(H)=\frac{(n-d)(n-d-1)}{2}$, $nt^{(1)}(H)=d(2d+1)$, $st^{(1)}(H)=d(2d-1)$.
@@ 6.5. Quasi dual hives
The goal of this section is to extend hexagon rotations to hives. As of now, hexagon rotations map one color map to another. To also change label maps, we need to relax the inequality constraints of Definition 5.2 leading to quasi hives of Definition 6.17.
@@
Let $n,d\geq 1$. Denote $E_{n,d}=\{\{u,v\}\in R_{n,d}^{2}:d(u,v)=1\}$ the set of edges of the discrete hexagon $R_{n,d}$. We define $H_{hex}(\lambda,\mu,\nu,N)$ as the set of maps $(C,L)$, $C:E_{n,d}\rightarrow\{0,1,3,m\}$ and $L:E_{n,d}\rightarrow\frac{1}{N}\mathbb{Z}$ satisfying conditions of Definition 5.2 restricted to edges $e\in E_{n,d}$.
By Lemma 6.9, any dual hive $H=(C,L)\in H(\lambda,\mu,\nu,N)$ with regular boundary conditions has the following property : every equilateral triangle of size $d$ anchored in a corner of $T_{n}$ has edges with color $C(e)=1$. Giving values of the label map $L$ restricted to any of these corner is thus equivalent via Proposition 5.8 to giving a classical Knutson-Tao puzzle. Since one of the boundaries of these puzzle is the empty partition by the regularity of boundaries, there is only one such puzzle and thus one such label map. Removing these fixed triangular corners, we have a bijection
@@
for regular boundaries. The inverse map is given by extending $C$ from the hexagon $R_{n,d}$ to the triangle $T_{n+d}$ by adding color $1$ edges in corners and completing the labels $L$ in the unique possible way in corner triangles. In this section, we now view hives in the hexagon $R_{n,d}$ and we will omit the subscript writing $H(\lambda,\mu,\nu,N)$ for notation convenience.
@@
Let $C:E_{n,d}\rightarrow\{0,1,3,m\}$ be a color map and $N\geq 1$. A quasi label map is a map $L:E_{n,d}\rightarrow\frac{1}{N}\mathbb{Z}$ satisfying the equality conditions of Definition 5.2, that is, equality condition on every triangular face and rigid lozenges with respect to the color map $C$. Denote $\widetilde{L}^{C}(\lambda,\mu,\nu,N)$ the set of such label maps.
@@
The difference with label maps of dual hives is that one does not impose the inequality constraints of Definition 5.2. Note that a dual two-colored hive is a quasi dual hive that is, $H(\lambda,\mu,\nu,N)\subset\widetilde{H}(\lambda,\mu,\nu,N)$.
@@
Let $H=(C,L)$ be a quasi dual hive and $h$ an ABC hexagon for its color map $C$. Then, the values of $L$ on $E_{h}^{\mathrm{o}}$ the interior edges of $h$ are uniquely determined by values of $L$ on boundary edges of $h$ and by the position of the rigid lozenge in $h$. Moreover, the values of $L$ on $E_{h}^{\mathrm{o}}$ are affine combinations of the values of $L$ on boundary edges of $h$.
@@
By the equality conditions on opposite sides of the rigid lozenge, one has $l_{12}=l_{1}$ and $l_{8}=l_{6}$. Let us show that the values $l_{7},l_{9},l_{10},l_{11}$ are uniquely determined in the region. Writing equality conditions on the four triangular faces gives
@@
Suppose that the former holds. The fourth equations of the system is redundant with the others and (6.4) so that $l_{7},l_{8},l_{9},l_{10},l_{11},l_{12}$ is solution to
@@
by setting $L^{\prime}(e)=L(e)$ for $e\in E_{n,d}\setminus E_{h}^{\mathrm{o}}$. Note that Lemma 6.18 ensures that the values of $L^{\prime}(e)$ for $e\in E_{h}^{\mathrm{o}}$ are uniquely determined by the new coloring $C^{\prime}$.
For notation convenience, we will call a hexagon rotation the application of a map of the form of Definition 6.19 for some ABC hexagon $h$ inside a quasi hive.
@@
Let $C,C^{\prime}$ be two color maps that differ by a hexagon rotation. The map $R[C\rightarrow C^{\prime}]:\ \widetilde{L}^{C}(\lambda,\mu,\nu,N)\rightarrow\widetilde{L}^{C^{\prime}}(\lambda,\mu,\nu,N)$ is an affine bijection with integer coefficients. Its inverse is $R[C^{\prime}\rightarrow C]$.
@@
For every edge $e$ not interior to $h$, one has $L^{\prime}(e)=L(e)$. In particular, $L^{\prime}(e)=L(e)$ for $e\in\partial h$. Using the same arguments as in the proof of Lemma 6.18, one shows that the values $\{L^{\prime}(e),e\in h^{\mathrm{o}}\}$ are integer combinations of values $\{L(e)=L^{\prime}(e),e\in\partial h\}$. ∎
Figure 29 shows an example of a hexagon rotation and the affine map $L\mapsto L^{\prime}$.
@@
$l^{\prime}_{9}$ $=l_{4}$ $l^{\prime}_{7}$ $=1-\frac{1}{N}-l_{6}-l^{\prime}_{12}=\frac{1}{N}+l_{3}+l_{5}-l_{6}$
$l^{\prime}_{11}$ $=l_{3}$ $l^{\prime}_{8}$ $=1-\frac{2}{N}-l_{1}-l^{\prime}_{7}=\frac{1}{N}+l_{3}+l_{5}-l_{6}$
$l^{\prime}_{12}$ $=1-\frac{2}{N}-l_{3}-l_{5}$ $l^{\prime}_{10}$ $=1-\frac{1}{N}-l_{4}-l^{\prime}_{3}.$

### v2

@@ 6.4. Color map reduction
Let $\{x^{(s)},1\leq s\leq d(k-d)\}$ be the vertices in $T_{k}$ ordered such that for $1\leq s\leq d(k-d)$, $s-1=s_{1}d+s_{2}$ with $s_{1}\geq 0$ and $10\leq s_{2}\leq d-$ 1,
$x^{(s)}=(d+s_{1})\xi+s_{2}\xi^{5}.$
Let $C$ be a regular color map on $E_{k}$. Let us show by induction on $s$ that using hexagon rotations, $C$ can be mapped to a regular color map $C^{(s)}$ such that the type 00 edges with origins $x^{(1)}+1,\dots,x^{(s)}+1$ are colored $m$. We first prove it for $s=1$. Notice that since $C$ is regular, there is a gash $g^{(1)}$ of type $2$ with center $x^{(1)}$. Applying Proposition 6.13 yields that using hexagon rotations, $C$ can be mapped to $C^{(1)}$ such that $g^{(1)}$ is adjacent to a configuration $(v)$ or $(vi)$. Since $C^{(1)}$ is regular, this configuration is necessarily $(v)$ which implies that the edge of type 00 with origin $x^{(1)}+1$ is colored $m$.
Assume that $C^{(s)}$ is a color map such that the type 00 edges with origins $x^{(1)}+1,\dots,x^{(s)}+1$ are colored $m$. Notice that there is a $01$ opening at $x^{(s+1)}$ that is, the edges $(x^{(s+1)},x^{(s+1)}+\xi)$ and $(x^{(s+1)},x^{(s+1)}+\xi^{5})$ are colored respectively 00 and $1$. A $01$ opening has only three possible completions showed in Figure 21.
@@
In the case of the first completion, the color map $C^{(s+1)}=C^{(s)}$ satisfies the desired conditions. In the case of the third and fourth completion, there is a gash $g^{(s+1)}$ of type $2$ and $1$ respectively with center $x^{(s+1)}+1$. Applying Proposition 6.13 and Remark 6.14 respectively shows that using hexagon rotations along the propagation path started from $g^{(s+1)}$, the latter is adjacent to a configuration $(v)$ or $(vi)$. Note that rotations in this propagation path do not affect edges colored $m$ with origins $x^{(1)}+1,\dots,x^{(s)}+1$ thanks to the above ordering. Using a hexagon rotation for the hexagon with center $x^{(s+1)}+1$ yields a color map $C^{(s+1)}$ such that the type 00 edges with origins $x^{(0)}+1,\dots,x^{(s+1)}+1$ are colored $m$ and ends the induction. Therefore, $C$ is equivalent up to hexagon rotations to the color map where type 00 edges with origins $x^{(0)}+1,\dots,x^{(d(k-d))}+1$ are colored $m$. From this configuration, there is only one possible color map to complete the rest of the hive $E_{k}$ which is the simple color map $C_{0}$. ∎
Since the number of pieces of each type is preserved under hexagon rotations, one derives the following enumerations (only the enumeration of edges colored $m$ will be used afterwards).
@@
Let $H$ be a dual hive with regular boundaries. Let $hc(H)$, respectively, $sc(H),$ be the number of $m$, respectively $3$ colored edges inside $H$. Then,
@@
Moreover, for $i\in\{0,1\}$, denote $nt^{(i)}(H)$, respectively $st^{(i)}(H)$, the number of direct, respectively reversed, triangular pieces of size $1$ with color $i$ on each side. Then,
@@
where $n(0)=n-d$ and $n(1)=2d$.
The detailed number of each type of triangles is thus
$nt^{(0)}(H)=\frac{(n-d)(n-d+1)}{2},\,st^{(0)}(H)=\frac{(n-d)(n-d-1)}{2},\,nt^{(1)}(H)=d(2d+1),\,st^{(1)}(H)=d(2d-1).$
@@ 6.5. Quasi dual hives
The goal of this section is to extend hexagon rotations to hives. As of now, hexagon rotations map one color map to another. To also change label maps, we need to relax the inequality constraints of Definition 5.1 leading to quasi hives of Definition 6.20. From this section to the end, we view regular dual hives of $T_{n+d}$ as in the discrete hexagon $R_{d,n}$, see Definitions 6.18, 6.19 and Figure 22 below.
@@
Let $n,d\geq 1$. Denote $E_{n,d}=\{\{u,v\}\in R_{d,n}^{2}:d(u,v)=1\}$ the set of edges of the discrete hexagon $R_{d,n}$. A hexagon dual hive is a pair of maps $(C,L)$, $C:E_{n,d}\rightarrow\{0,1,3,m\}$ and $L:E_{n,d}\rightarrow\frac{1}{N}\mathbb{Z}$ such that $NL(.)$ satisfies the conditions of Definition 5.1 restricted to edges $e\in E_{n,d}$.
Definition 6.19 (Boundary value of a hexagonal dual hives).
Define the following subsets of $E_{n,d}$ for $l\in\{0,1,2\}$.
$\partial^{(l,l)}E_{n,d}$ =$=\{e\in E_{n,d}:e\text{ is of type }l\text{ and }e_{l-1}=0\}$
$\partial^{(l,l+1)}E_{n,d}$ =$=\{e\in E_{n,d}:e\text{ is of type }l\text{ and }e_{l-1}=n\}.$
The boundary value $[(c^{(0,0)},c^{(1,1)},c^{(2,2)},c^{(0,1)},c^{(1,2)},c^{(2,0)}),(\ell^{(0,0)},\ell^{(1,1)},\ell^{(2,2)},\ell^{(0,1)},\ell^{(1,2)},\ell^{(2,0)})]$ of a hexagonal dual hive is the restriction of $(C,L)$ to $\partial E_{n,d}$ where
  • $c^{(l,l)}\in\{0,1,3,m\}^{n-d}$ (resp. $c^{(l,l+1)}\in\{0,1,3,m\}^{d}$) is the restriction of $C$ to $\partial^{(l,l)}E_{n,d}$ (resp. $\partial^{(l,l+1)}E_{n,d}$ ),
  • $\ell^{(l,l)}\in(\frac{1}{N}\mathbb{N})^{d}$ (resp. $\ell^{(l,l+1)}\in(\frac{1}{N}\mathbb{N})^{d}$) is the restriction of $L(.)$ to $\partial^{(l,l)}E_{n,d}$ (resp. $\partial^{(l,l+1)}E_{n,d}$).
For $(c,l)=(c^{(0,0)},c^{(1,1)},c^{(2,2)},c^{(0,1)},c^{(1,2)},c^{(2,0)},\ell^{(0,0)},\ell^{(1,1)},\ell^{(2,2)},\ell^{(0,1)},\ell^{(1,2)},\ell^{(2,0)})\in\{0,1\}^{3(n-d)}\times\{0,1\}^{3d}\times(\frac{1}{N}\mathbb{N})^{3(n-d)}\times(\frac{1}{N}\mathbb{N})^{3d}$, we denote by $\mathcal{H}_{hex}(c,l)$ the set of hexagonal dual hives with boundary value $(c,l)$.
Let $(\lambda,\mu,\nu)$ be partitions of length $n$ with first part smaller than $N-n$ such that $|\lambda|+|\mu|=|\nu|+Nd$ and such that $\min(\lambda_{n},\mu_{n},N-n-\nu_{1})\geq d-1$. Then, the associated boundary labels $(l^{(0)},l^{(1)},l^{(2)})$ defined in Definition 5.1 on $T_{n+d}$ are given by
$l^{(0)}$ $=(0,\dots,d-1,N-n-\nu_{1},N-n-\nu_{2}+1,\dots,N-\nu_{n-d}-d-1,N-\nu_{n-d+1}-d,\dots,N-\nu_{n}-1)$
$l^{(1)}$ $=(0,\dots,d-1,\mu_{n},\mu_{n-1}+1,\dots,\mu_{d+1}+n-d-1,\mu_{d}+n-d,\dots,\mu_{1}+n-1)$
$l^{(2)}$ $=(0,\dots,d-1,\lambda_{n},\lambda_{n-1}+1,\dots,\lambda_{d+1}+n-d-1,\lambda_{d}+n-d,\dots,\lambda_{1}+n-1)$
so that the associated boundary colors $(c^{(0)},c^{(1)},c^{(2)})$ are regular in the sense of Definition 6.1 which means that on each boundary of $T_{n+d}$, the $d$ first and last edges are colored $1$ and the remaining $n-d$ edges are colored 00. By Lemma 6.10, any dual hive $H=(C,L)\in H(\lambda,\mu,\nu,N)$ with regular boundary conditions has every of its equilateral triangles of size $d$ anchored in a corner of $T_{n+d}$ colored $1$. Each of these triangular regions have one boundary with labels equal to $(0,\dots,d-1)$. The corresponding region in the puzzle yields a unique position of the corresponding triangular pieces with edge colors $(1,1,1)$ and this unique configuration gives labels to the third boundary edge of the region. Let $\tilde{l}^{(0)},\tilde{l}^{(1)},\tilde{l}^{(2)}$ be the respective labels of edges in $\partial^{(0,1)}E_{n,d},\partial^{(1,2)}E_{n,d},\partial^{(2,0)}E_{n,d}$. Then, reading decreasingly with respect to edges heights $h(e)$, these labels are given by the following (see Figure 22 below),
(6.4) $\tilde{l}^{(0)}$ $=(N-n-\lambda_{d}+d-1,N-n-\lambda_{d+1}+d,\dots,N-n-\lambda_{1}),$
(6.5) $\tilde{l}^{(1)}$ $=(N-n-\mu_{d}+d-1,N-n-\mu_{d+1}+d,\dots,N-n-\mu_{1}),$
(6.6) $\tilde{l}^{(2)}$ $=(\nu_{n-d+1}+d-1,\dots,\nu_{n-1}+1,\nu_{n}).$
Figure 22. The restriction of a regular dual hive on $E_{n+d}$ to the hexagon $E_{n,d}$ and induced boundary value for the label map. Blue labels are determined by the unique label map on each triangular corner. Multiplying edge labels by $\frac{1}{N}$ yields boundary conditions in $H_{hex}(\lambda,\mu,\nu,N)$.
In this section, we now view regular dual hives $H=(C,L)$ on $T_{n+d}$ as hexagonal dual hives by restricting $C$ and $L$ to $E_{n,d}$. By the above, the boundary conditions on $E_{n,d}$ are given by
(6.7) $c^{(l,l)}=(0,\dots,0),\ c^{(l,l+1)}=(1,\dots,1),\ \ell^{(l,l)}=\frac{1}{N}L_{|\partial^{(l,l)}E_{n,d}},\ \ell^{(l,l+1)}=\frac{1}{N}\tilde{l}^{(l)}.$
for $l\in\{0,1,2\}$. We write $H_{hex}(\lambda,\mu,\nu,N)$ for the set of hexagonal dual hives having boundary colors and labels given by (6.7) coming from restriction of dual hives on $T_{n+d}$. This restriction
@@
is a bijection where the inverse map is given by extending $C$ from the hexagon $R_{d,n}$ to the triangle $T_{n+d}$ setting $C(e)=1$ for $e\in E_{n+d}\setminus E_{n,d}$ and completing the labels $L$ in the unique possible way in corner triangles. In this section, we now view dual hives in the hexagon $R_{d,n}$ and we will omit the subscript writing $H(\lambda,\mu,\nu,N)$ for notation convenience.
@@
Let $C:E_{n,d}\rightarrow\{0,1,3,m\}$ be a color map and $N\geq 1$. A quasi label map is a map $L:E_{n,d}\rightarrow\frac{1}{N}\mathbb{Z}$ such that $NL(.)$ satisfying the equality conditions of Definition 5.1, that is, equality condition on every triangular face and rigid lozenges with respect to the color map $C$, and boundary values on $\partial E_{n,d}$ given by the corresponding two-colored dual hives. Denote $\widetilde{L}^{C}(\lambda,\mu,\nu,N)$ the set of such label maps.
@@
The difference with label maps of dual hives is that one does not impose the inequality constraints of Definition 5.1 inside the hexagonal region $E_{n,d}$. Note that dual two-colored hives are in particular quasi dual hives that is, $H(\lambda,\mu,\nu,N)\subset\widetilde{H}(\lambda,\mu,\nu,N)$.
@@
Let $H=(C,L)$ be a quasi dual hive and $h$ an ABC hexagon for its color map $C$. Then, the values of $L$ on $E_{h}^{\mathrm{o}}$, the set of interior edges of $h$, are uniquely determined by the values of $L$ on boundary edges of $h$ and by the position of the rigid lozenge in $h$. Moreover, the values of $L$ on $E_{h}^{\mathrm{o}}$ are affine combinations of the values of $L$ on boundary edges of $h$.
@@
By the equality conditions on opposite sides of the rigid lozenge, one has $l_{12}=l_{1}$ and $l_{8}=l_{6}$. Let us show that the values $l_{7},l_{9},l_{10},l_{11}$ are uniquely determined in the region. Writing equality conditions on the five triangular faces gives
@@
Suppose that the former holds. The fourth equations of the system is redundant with the others and (6.9) so that $l_{7},l_{8},l_{9},l_{10},l_{11},l_{12}$ is solution to the invertible system
@@
The labels on the inner edges are thus uniquely determined as linear combinations of the labels on the outer edges.
@@
by setting $L^{\prime}(e)=L(e)$ for $e\in E_{n,d}\setminus E_{h}^{\mathrm{o}}$ and extending $L^{\prime}$ to $E_{h}^{\mathrm{o}}=E_{h^{\prime}}^{\mathrm{o}}$ by Lemma 6.21.
For notation convenience, we will call a hexagon rotation the image of a map of Definition 6.22 for some ABC hexagon $h$ inside a quasi hive.
@@
Let $C,C^{\prime}$ be two color maps that differ by a hexagon rotation. The map $R[C\rightarrow C^{\prime}]:\ \widetilde{L}^{C}(\lambda,\mu,\nu,N)\rightarrow\widetilde{L}^{C^{\prime}}(\lambda,\mu,\nu,N)$ is an affine bijection with coefficients in $\mathbb{Z}[\lambda_{i},\mu_{i},\nu_{i},1/N]$ and whose inverse is $R[C^{\prime}\rightarrow C]$.
@@
For every edge $e$ not interior to $h$, one has $L^{\prime}(e)=L(e)$. In particular, $L^{\prime}(e)=L(e)$ for $e\in\partial h$. Since the matrix of (6.10) has integer coefficients and is lower triangular, the values $\{L^{\prime}(e),e\in E_{h}^{\mathrm{o}}\}$ are integer combinations of the values $\{L(e)=L^{\prime}(e),e\in\partial h\}$ and $\frac{1}{N}\mathbb{Z}$. ∎
Figure 24 shows an example of a hexagon rotation and the corresponding affine map $L\mapsto L^{\prime}$.
@@
$l^{\prime}_{9}$ $=l_{4},$ $l^{\prime}_{7}$ $=1-\frac{1}{N}-l_{6}-l^{\prime}_{12}=\frac{1}{N}+l_{3}+l_{5}-l_{6},$
$l^{\prime}_{11}$ $=l_{3},$ $l^{\prime}_{8}$ $=1-\frac{2}{N}-l_{1}-l^{\prime}_{7}=\frac{1}{N}+l_{3}+l_{5}-l_{6},$
$l^{\prime}_{12}$ $=1-\frac{2}{N}-l_{3}-l_{5},$ $l^{\prime}_{10}$ $=1-\frac{1}{N}-l_{4}-l^{\prime}_{3}.$